### v1

@@ 3 An infinite $\beta$-FPUT chain with an isotopic defect
Finally, consider the case of a light defect $10<m<$ 1. As it is shown in Shishkina2023cmat , in the corresponding linear system, the localized component ${v}_{n}^{\mathrm{stop}}$ co-exists with the anti-localized one ${v}_{n}^{\mathrm{pass}}$. We also know that adding nonlinearity does not destroy an impurity-induced localized mode in the case of the chain Kivshar1991 . In Fig. 4, where the particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time, one can see that in the nonlinear system we also observe a similar wave pattern, which corresponds to the superposition of localized and anti-localized components. In Fig. 5, the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (3.25) is also displayed. In both cases (the nonlinear and the linear ones), a non-vanishing impurity induced localized oscillation is observed.
@@
Figure 2: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a heavy isotopic defect. The anti-localization, i.e., zeroing of the perturbation amplitude near $n=0$, is clearly observed for both linear and nonlinear systems with the defect and not observed for systems without the defect
Figure 3: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solutions for a system with a heavy isotopic defect; (b) the solutions for the corresponding uniform system. In both sub-plots, the order of decay for the nonlinear system can be estimated as the same as it is observed in the corresponding linear system
@@
Figure 5: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solution for a system with a light isotopic defect; (b) the solution for the corresponding uniform system. For both linear and nonlinear systems, a non-vanishing localized oscillation is observed
4 An infinite $\beta$-FPUT chain with a mass-spring inclusion
As we can see from the previous Sect. 3, the anti-localization exists in linear and nonlinear systems with an isotopic defect. The anti-localization is observed for both cases $m>1$ and $m<1$, though, in the latter case, the anti-localization co-exists with the localization. The boundary in the system parameter space between these two cases corresponds to the uniform system, where there is no anti-localization. In previous study Shishkina2023jsv , we have shown that besides the uniform system, the anti-localization disappears in linear non-uniform systems with parameters, which correspond to such a boundary separating domains in the system parameter space where the localization is observed and not observed. The linear uniform system is just a particular case of such a system with the boundary parameter values. In this section, we want to check if the anti-localization disappears for non-uniform nonlinear systems with such boundary values of their parameters. To achieve this, we consider the chain with parameters satisfying Eq. (3.3), equipped with a mass-spring inclusion for which Eq. (3.2) and
$K\neq 1$ (4.1)
are fulfilled. The defect is subjected to a pulse loading at the position defined by Eq. (3.5). In the corresponding linear system with $B=\beta=0$ the impurity-induced localization is observed if and only if the following inequality is fulfilled Montroll1955 ; Yu2019 ; Shishkina2023asm :
$K>K_{0},$ (4.2)
$K_{0}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\frac{2m}{1+m}.$ (4.3)
Inequality (4.2) defines a domain in the problem parameter space, which we call the localization domain. Now consider a nonlinear system with a fixed value $B=\beta$, and fixed $m$ as we have taken in Sect. 3. Now we try to determine the boundary value of the parameter $K$, which separates nonlinear systems where the impurity-induced localization is observed or not observed. The numeric experiment shows that such a value is very close to the value $K=K_{0}$ defined by Eq. (4.3) for the linear systems, see Figs. 68.
@@ 4 An infinite $\beta$-FPUT chain with a mass-spring inclusion
In Fig. 6 the numerical solution $v_{0}(t)$ versus the time $t$ for $\beta$-FPUT chain with a mass-spring defect is presented. The value of the linear spring stiffness is taken equal to the boundary value $K_{0}$. For reference, the same particle velocity in the corresponding linear system with the same defect, as well as the exact solution (2.9) for the uniform linear chain, are also presented. [^5] Unfortunately, we do not have here the reference asymptotic non-stationary solution for the corresponding linear system. The order of decay for oscillation in both nonlinear systems with and without the defect can be estimated as $1/\sqrt{t}$ as it is observed according to Eqs. (3.17), (3.26) in the uniform linear chain. In Fig. 7, the particle velocities $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ are shown at a fixed value of time for the same three systems as in Fig. 6. The absence of the anti-localization in the system with the defect for which $K=K_{0}$ is clearly recognizable.
In Fig. 8 the same numerical solution $v_{0}(t)$ for $\beta$-FPUT chain with a mass-spring defect with $K=K_{0}$, which is already shown in Fig. 6, is compared with analogous solutions with $K=0.9K_{0}$ and $K=1.1K_{0}$. We can see that the solution corresponding to $K=0.9K_{0}$ vanishes faster than one taken for $K=K_{0}$. The corresponding order of decay is expected to be $1/\sqrt{t^{3}}$, see Sect. 3. On the other hand, the solution corresponding to $K=1.1K_{0}$ is a non-vanishing one, i.e., an impurity induced localized oscillation is observed.
@@
5 A semi-infinite $\beta$-FPUT chain
Assuming that $K=0$, $B=0$, we split an infinite chain into two uncoupled semi-infinite chains. Thus, a semi-infinite chain loaded at the end $n=N=1$ is a particular case of a chain with a defect. The exact solution is Liazhkov2025
$v_{n}=J_{2|n-1|}(2t)+J_{2n}(2t).$ (5.1)
The anti-localization in such a linear chain ($\beta=0$) was demonstrated in study Gavrilov2023DD , where the asymptotic solution for $v_{n}$ in the form of the right-hand side of Eq. (3.17) wherein
${A}(w)=\frac{2w}{\sqrt{\pi}\sqrt[4]{1-w^{2}}},$ (5.2)
$\psi=\arctan\frac{\sqrt{1-w^{2}}}{w},$ (5.3)
$w=\frac{|n-1|}{t}$ (5.4)
was obtained. The expression for $\phi$ coincides with (3.22). Clearly, $A(0)=0$, which implies the anti-localization effect. For the end particle, one has Gavrilov2023DD :
$v_{1}=\frac{\sin(2t-\frac{\pi}{4})}{\sqrt{\pi}t^{3/2}}+o(t^{-3/2}).$ (5.5)
This effect was discovered for the first time, as far as we know now [^6] We are grateful to S.D. Liazhkov, much earlier in Jackson1978 and explained by specific peculiarities of linear systems. Let us check if the anti-localization exists in a semi-infinite $\beta$-FPUT chain.
The numerical solution, which corresponds to a nonlinear chain with $\beta\neq 0$ is presented in Fig. 9. For reference, the wave-field in the corresponding linear semi-infinite chain given by the exact solution (5.1) is also displayed. In Fig. 9a the particle velocities $v_{n}$ versus the spatial variable $n$ are presented. In Fig. 9b the particle velocity of the end particle $v_{1}(t)$ versus time is shown. One can see that the anti-localization is clearly visible for both linear and nonlinear systems.
Remark 5.1
In Fig. 9a one can also observe a soliton-like wave propagating faster than the leading wave-front in the linear system. The solitons in a $\beta$-FPUT chain are discussed, e.g., in Vainchtein2022 .
@@ 5 A semi-infinite $\beta$-FPUT chain
6 A nonlinear spring in an infinite linear chain
@@
$m=1,\quad K=1,\quad\beta=0.$ (6.1)
@@
$B_{L}=0,\quad B_{R}>0.$ (6.2)
In Fig. 10 we present the spatial distribution of the particle velocities in the corresponding numeric experiment with pulse loading applied at the particle with number $n=N=0$ just to the left from the nonlinear spring. Sub-plots (a) & (b) correspond to the choices $B_{R}=4.0$ and $B_{R}=10.0$, respectively. For reference, the particle velocities in the corresponding linear system with $B_{L}=B_{R}=0$ given by Eq. (2.9) are also displayed in each sub-plot. One can see that the wave pattern observed in Fig. 10b is a bit similar to the one presented in Fig. 4, where the anti-localization co-exists with the impurity-induced localized oscillation. The clear difference is an expected asymmetry of the wave patterns in Fig. 10. In Fig. 10a we also observe weakening of the propagating component near the defect but not zeroing. Increasing the coefficient $B_{R}$ one can see that the amplitude of the propagating component near the defect decreases. Thus, we observe a weakening of the propagating component but, apparently, not an asymptotic one. The observed phenomenon can be called the incomplete anti-localization. Here we cannot estimate, even qualitatively, the order of decay for the anti-localized component due to the presence of the localized mode.
@@
A possibility to eliminate the localized oscillation is to consider a pulse loading applied at some distance $N>0$ outside the defect. For such a case, in study Gavrilov2024 , we have shown that in the linear chain the anti-localization is observed in some one-sided neighbourhood just behind the defect, which expands with time. The amplitude of the localized oscillation in the linear system exponentially decreases with $N$; therefore this oscillation is not observable for large enough $N$. The order of decay for the particle velocities in such a shadow zone is the same as it is observed when the loading is applied at a defect, i.e., $t^{-3/2}$. In Fig. 11, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ in an infinite linear chain, with an inclusion in the form of one strongly nonlinear spring, is shown for a fixed value of time. Here, we take an extremely large value $B_{R}=1000$ to obtain a clearly visible effect. For reference, the wave-field in the uniform linear chain given by Eq. (2.9), is also displayed. One can compare Fig. 11 with Fig. 1 in Shishkina2023cmat , where the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for the analogous problem for the linear chain with an isotopic defect. The plots are very similar, and the shadow zones behind the defect are clearly recognizable in both cases.
@@
In Fig. 12 the numerically found solution for the particle velocity $v_{-1}(t)$ versus the time $t$ is shown for the same system. The choice $n=-1$ corresponds to the first particle behind the defect, which is not connected to the defective spring. One can see that the estimate for the order of the decay should be taken as $t^{-1/2}$, not $t^{-3/2}$ as usual when the (complete) anti-localization is observed. Thus, we really observe a weakening of the oscillation behind the defect, but not an asymptotic one. The leading order term of the large time asymptotics for the oscillation amplitude is relatively small, but not zero.
@@
7 Conclusion
In the paper, we have presented the numerical results for investigation of several nonlinear problems concerning the quasi-waves propagation in an infinitely long $\beta$-FPUT chain with a defect. Though we have demonstrated the results, which correspond to a specific choice of the problem parameters, in our work, we have considered a wide band of them. This allows us to claim that the phenomenon of anti-localization of non-stationary waves is not somehow related to the linearity of the system under consideration. Anti-localization is natural for both linear and nonlinear systems with defects.
@@ 7 Conclusion
    An anti-localized propagating component of a wave-field may co-exist with an impurity-induced localized wave; see, e.g., Fig. 4, which corresponds to a light isotopic defect. The localized mode forms a peak of the amplitude near a defect, but this peak is observed inside a zone, where the oscillation amplitude is zero.
@@
    The anti-localization is not observed for systems with parameters at the boundary of the domain in the problem parameter space where the wave localization occurs; see, e.g., Fig. 7, which corresponds to a mass-spring defect. Thus, we have shown that the absence of the anti-localization in a uniform system is not related to the uniformity itself. A uniform system is just the particular case of a system with such boundary parameters.
@@
This was shown previously in Shishkina2023jsv for a continuum linear system in an analytic way. In this paper, we have demonstrated numerically an analogous behaviour of nonlinear systems with finite nonlinearity.
@@
In Sect. 6, we have demonstrated that embedding the point strong nonlinearities into a linear chain can also lead to a weakening of the propagating component of the wave-field in a neighbourhood of such a defect. Though this effect is quite similar to one which we call the anti-localization, in the latter case, we observe weakening but not zeroing of the leading-order term. This can be called “the incomplete anti-localization”.
@@
Note that introducing nonlinearity into a linear system possessing a localized mode with the corresponding frequency inside a stop-band, in some cases, can lead to destroying the mode. This happens at least in the case when the stop-band correspond to smaller values of frequency compared with the pass-band kaplunov2008example ; indeitsev2004localization due to the emergence of super-linears, whose frequencies are multiples of the localized mode frequency. These frequencies can leave the stop-band yielding the leakage of energy of the localized mode. In this paper, we consider systems with a stop-band corresponding to higher values of the frequency compared with a pass-band. This situation is quite usual for discrete systems with zero on-site potential. The anti-localization is the phenomenon, which is closely related to the localization; thus, in nonlinear systems where a stop-band corresponds to small values of the frequency (this is typical for the continuous case) the anti-localization requires an additional investigation.

### v2

@@ III An infinite $\beta$-FPUT chain with an isotopic defect
Finally, consider the case of a light defect $0<m<1$. As it is shown in Shishkina and Gavrilov (2023a), in the corresponding linear system, the localized component ${v}_{n}^{\mathrm{stop}}$ co-exists with the anti-localized one ${v}_{n}^{\mathrm{pass}}$. We also know that adding nonlinearity does not destroy an impurity-induced localized mode in the case of the chain Kivshar (1991). In Fig. 4, where the particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time, one can see that in the nonlinear system we also observe a similar wave pattern, which corresponds to the superposition of localized and anti-localized components. In Fig. 5, the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (43) is also displayed. In both cases (the nonlinear and the linear ones), a non-vanishing impurity induced localized oscillation is observed. The frequency shift observed for the nonlinear system confirms the nonlinear character of the localized oscillation.
@@
Figure 2: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a heavy isotopic defect. The anti-localization, i.e., zeroing of the perturbation amplitude near $n=0$, is clearly observed for both linear and nonlinear systems with the defect and not observed for systems without the defect. The difference between nonlinear and linear solutions for the system with the defect is clearly recognizable everywhere excepting the anti-localization zone.
Figure 3: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solutions for a system with a heavy isotopic defect; (b) the solutions for the corresponding uniform system. In both sub-plots, the order of decay for the nonlinear system can be estimated as the same as it is observed in the corresponding linear system. Oscillation asymmetry is clearly recognizable for the nonlinear system with the defect.
@@
Figure 5: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain in the case of a light isotopic defect. For both linear and nonlinear systems, a non-vanishing localized oscillation is observed. The frequency shift observed for the nonlinear system confirms the nonlinear character of the oscillation
IV An infinite $\beta$-FPUT chain with a mass-spring inclusion
As we can see from the previous Sect. III, the anti-localization exists in linear and nonlinear systems with an isotopic defect. The anti-localization is observed for both cases $m>1$ and $m<1$, though, in the latter case, the anti-localization co-exists with the localization. The boundary in the system parameter space between these two cases corresponds to the uniform system, where there is no anti-localization. In previous study Shishkina et al. (2023), we have shown that besides the uniform system, the anti-localization disappears in linear non-uniform systems with parameters, which correspond to such a boundary separating domains in the system parameter space where the localization is observed and not observed. The linear uniform system is just a particular case of such a system with the boundary parameter values. In this section, we want to check if the anti-localization disappears for non-uniform nonlinear systems with such boundary values of their parameters. To achieve this, we consider the chain with parameters satisfying Eq. (12), equipped with a mass-spring inclusion for which Eq. (11) and
$K\neq 1$ (15)
are fulfilled. The defect is subjected to a pulse loading at the position defined by Eq. (14). In the corresponding linear system with $B=\beta=0$ the impurity-induced localization is observed if and only if the following inequality is fulfilled Montroll and Potts (1955); Yu (2019); Shishkina and Gavrilov (2023b):
$K>K_{0},$ (16)
$K_{0}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\frac{2m}{1+m}.$ (17)
Inequality (16) defines a domain in the problem parameter space, which we call the localization domain. Now consider a nonlinear system with a fixed value $B=\beta$, and fixed $m$ as we have taken in Sect. III. Now we try to determine the boundary value of the parameter $K$, which separates nonlinear systems where the impurity-induced localization is observed or not observed. The numeric experiment shows that such a value is very close to the value $K=K_{0}$ defined by Eq. (17) for the linear systems, see Figs. 68.
@@ IV An infinite $\beta$-FPUT chain with a mass-spring inclusion
In Fig. 6 the numerical solution $v_{0}(t)$ versus the time $t$ for $\beta$-FPUT chain with a mass-spring defect is presented. The value of the linear spring stiffness is taken equal to the boundary value $K_{0}$. For reference, the same particle velocity in the corresponding linear system with the same defect, as well as the exact solution (9) for the uniform linear chain, are also presented  [^5] Unfortunately, we do not have here the reference asymptotic non-stationary solution for the corresponding linear system.. The order of decay for oscillation in both nonlinear systems with and without the defect can be estimated as $1/\sqrt{t}$ as it is observed according to Eqs. (35), (44) in the uniform linear chain. In Fig. 7, the particle velocities $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ are shown at a fixed value of time for the same three systems as in Fig. 6. The absence of the anti-localization in the system with the defect for which $K=K_{0}$ is clearly recognizable.
In Fig. 8 the same numerical solution $v_{0}(t)$ for $\beta$-FPUT chain with a mass-spring defect with $K=K_{0}$, which is already shown in Fig. 6, is compared with analogous solutions with $K=0.9K_{0}$ and $K=1.1K_{0}$. We can see that the solution corresponding to $K=0.9K_{0}$ vanishes faster than one taken for $K=K_{0}$. The corresponding order of decay is expected to be $1/\sqrt{t^{3}}$, see Sect. III. On the other hand, the solution corresponding to $K=1.1K_{0}$ is a non-vanishing one, i.e., an impurity induced localized oscillation is observed.
@@
V A semi-infinite $\beta$-FPUT chain
Assuming that $K=0$, $B=0$, we split an infinite chain into two uncoupled semi-infinite chains. Thus, a semi-infinite chain loaded at the end $n=N=1$ is a particular case of a chain with a defect.
For the corresponding linear semi-infinite system, the effect that we call the anti-localization was first time discovered, as far as we know  [^6] We are grateful to S.D. Liazhkov., much earlier in Jackson (1978) and explained by specific peculiarities of linear systems. The basic formulas describing anti-localized wave-field in the linear case the reader can find in Appendix C. Let us check if the anti-localization exists in a semi-infinite $\beta$-FPUT chain.
The numerical solution, which corresponds to a nonlinear chain with $\beta\neq 0$ is presented in Fig. 9. For reference, the wave-field in the corresponding linear semi-infinite chain given by the exact solution (45) is also displayed. In Fig. 9a the particle velocities $v_{n}$ versus the spatial variable $n$ are presented. In Fig. 9b the particle velocity of the end particle $v_{1}(t)$ versus time is shown. One can see that the anti-localization is clearly visible for both linear and nonlinear systems.
Remark 1.
In Fig. 9a one can also observe a soliton-like wave propagating faster than the leading wave-front in the linear system. The solitons in a $\beta$-FPUT chain are discussed, e.g., in Vainchtein (2022).
@@ V A semi-infinite $\beta$-FPUT chain
VI A nonlinear spring in an infinite linear chain
@@
$m=1,\quad K=1,\quad\beta=0.$ (18)
@@
$B_{L}=0,\quad B_{R}>0.$ (19)
In Fig. 10 we present the spatial distribution of the particle velocities in the corresponding numeric experiment with pulse loading applied at the particle with number $n=N=0$ just to the left from the nonlinear spring. Sub-plots (a) & (b) correspond to the choices $B_{R}=4.0$ and $B_{R}=10.0$, respectively. For reference, the particle velocities in the corresponding linear system with $B_{L}=B_{R}=0$ given by Eq. (9) are also displayed in each sub-plot. One can see that the wave pattern observed in Fig. 10b is a bit similar to the one presented in Fig. 4, where the anti-localization co-exists with the impurity-induced localized oscillation. The clear difference is an expected asymmetry of the wave patterns in Fig. 10. In Fig. 10a we also observe weakening of the propagating component near the defect but not zeroing. Increasing the coefficient $B_{R}$ one can see that the amplitude of the propagating component near the defect decreases. Thus, we observe a weakening of the propagating component but, apparently, not an asymptotic one. The observed phenomenon can be called the incomplete anti-localization. Here we cannot estimate, even qualitatively, the order of decay for the anti-localized component due to the presence of the localized mode.
@@
A possibility to eliminate the localized oscillation is to consider a pulse loading applied at some distance $N>0$ outside the defect. For such a case, in study Gavrilov and Shishkina (2024), we have shown that in the linear chain the anti-localization is observed in some one-sided neighborhood just behind the defect, which expands with time. The amplitude of the localized oscillation in the linear system exponentially decreases with $N$; therefore this oscillation is not observable for large enough $N$. The order of decay for the particle velocities in such a shadow zone is the same as it is observed when the loading is applied at a defect, i.e., $t^{-3/2}$. In Fig. 11, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ in an infinite linear chain, with an inclusion in the form of one strongly nonlinear spring, is shown for a fixed value of time. Here, we take an extremely large value $B_{R}=1000$ to obtain a clearly visible effect. For reference, the wave-field in the uniform linear chain given by Eq. (9), is also displayed. One can compare Fig. 11 with Fig. 1 in Shishkina and Gavrilov (2023a), where the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for the analogous problem for the linear chain with an isotopic defect. The plots are very similar, and the shadow zones behind the defect are clearly recognizable in both cases.
@@
In Fig. 12 the numerically found solution for the particle velocity $v_{-1}(t)$ versus the time $t$ is shown for the same system. The choice n=1n=-1 corresponds to the first particle behind the defect, which is not connected to the defective spring. One can see that the estimate for the order of the decay should be taken as $t^{-1/2}$, not $t^{-3/2}$ as usual when the (complete) anti-localization is observed. Thus, we really observe a weakening of the oscillation behind the defect, but not an asymptotic one. The leading order term of the large time asymptotics for the oscillation amplitude is relatively small, but not zero.
@@
VII Conclusion
In the paper, we have presented the numerical results for the investigation of several strongly nonlinear problems concerning the quasi-waves propagation in an infinitely long $\beta$-FPUT chain with a defect. Though we have demonstrated the results, which correspond to a specific choice of the problem parameters, in our work, we have considered a wide band of them. This allows us to claim that the phenomenon of anti-localization of non-stationary waves is not somehow related to the linearity of the system under consideration. Anti-localization is natural for both linear and nonlinear systems with defects.
@@ VII Conclusion
    An anti-localized propagating component of a wave-field may co-exist with an impurity-induced localized wave; see, e.g., Fig. 4, which corresponds to a light isotopic defect. The localized mode forms a peak of the amplitude near a defect, but this peak is observed inside a zone, where the oscillation amplitude is almost zero.
@@
    The anti-localization is not observed for systems with parameters at the boundary of the localization domain  [^7] A domain in the problem parameter space where the wave localization occurs.; see, e.g., Fig. 7, which corresponds to a mass-spring defect. Thus, we have shown that the absence of the anti-localization in a uniform system is not related to the uniformity itself. A uniform system is just the particular case of a system with such boundary parameters.
@@
These results were shown previously in Shishkina et al. (2023) for a continuum linear system in an analytic way. In this paper, we have demonstrated numerically an analogous behavior of strongly nonlinear systems.
@@
In Sect. VI, we have demonstrated that embedding of the point strong nonlinearities into a linear chain can also lead to a weakening of the propagating component of the wave-field in a neighborhood of such a defect. Though this effect is quite similar to one which we call the anti-localization, in the latter case, we observe weakening but not zeroing of the leading-order term. This can be called ‘‘the incomplete anti-localization’’.
Although the conclusions in the current paper are very similar to ones previously obtained for the linear case, we claim that they cannot be firmly deduced from any linear consideration. The observations that support this point of view are as follows:
  • The considered nonlinearity clearly quantitatively alters the solution at least everywhere excepting the neighborhood of the defect; see, e.g., Figs. 2, 9a;
  • The considered nonlinearity clearly quantitatively alters solution at the defect for the finite (not very small and not very large) values of time; see, e.g., Figs. 3a, 9b;
  • In the case of a heavy isotopic defect, nonlinearity qualitatively alters the solution at the defect for finite and large times. Namely, oscillation asymmetry is observed; see Fig. 3a. This effect is impossible Shishkina and Gavrilov (2023a) in the linear system.
  • Anti-localization can be observed in a neighborhood of a nonlinear impurity-induced localized mode; see, e.g., Fig. 4;
  • The phenomenon of the incomplete anti-localization discussed in Sect. VI is observed for a nonlinear system, which is uniform in the linear approximation;
  • The anti-localization in a linear semi-infinite chain has been demonstrated in the strongly nonlinear case; see Fig. 9; although before (Jackson (1978), p. 137) this effect was explained as a specific peculiarity of a linear system;
In our opinion, a nonlinearity, of course, could destroy the anti-localization as suggested in Jackson (1978). Indeed, in domains where localized waves are not observed, the linear equations are applicable for large times. However, the initial and boundary conditions for these equations are completely different from ones that we consider investigating the anti-localization in linear systems. Nonlinear localized waves of a finite amplitude (impurity-induced ones and solitary waves at the leading wave-fronts), in principle, could destroy the anti-localization. These localized structures are the sources, from which the perturbations propagate into a neighborhood of the defect. In Gavrilov and Shishkina (2024) we have shown that in the linear case a source outside the defect destroys anti-localization inside a zone between the source and the defect. The essentially nonlinear stage of the motion, which corresponds to finite values of time, also can affect defect’s oscillation at large times, see Fig. 3a. Nevertheless, the investigation in this paper shows that the anti-localization survives in nonlinear systems. Thus, the anti-localization can also be observed in weakly nonlinear systems for very large times, i.e., in the thermodynamic limit.
@@
Note that introducing nonlinearity into a linear system possessing a localized mode with the corresponding frequency inside a stop-band, in some cases, can lead to destroying the mode. This happens at least in the case when the stop-band correspond to smaller values of frequency compared with the pass-band Kaplunov and Nolde (2008); Indeitsev and Osipova (2004) due to the emergence of super-harmonics, whose frequencies are multiples of the localized mode frequency. These frequencies can leave the stop-band yielding the leakage of energy of the localized mode. In this paper, we consider systems with a stop-band corresponding to higher values of the frequency compared with a pass-band. This situation is quite usual for discrete systems with zero on-site potential. The anti-localization is the phenomenon, which is closely related to the localization; thus, in nonlinear systems where a stop-band corresponds to small values of the frequency (this is typical for the continuous case) the anti-localization requires an additional investigation.